Beam Inc. bonds are trading today for a price of $800.26 The bond currently has 22 years until maturity and has a yield to maturity of 7.36​%.

The bond pays annual coupons and the next coupon is due in one year. What is the coupon rate of the​ bond?

The coupon rate of the bond is __ ​%. (Round to one decimal​ place.)
Answer #1

price of coupon = Coupon payment per period * [1-(1+i)^-n]/i + par value/(1+i)^n

i = interest rate per period

n = number of periods

=>

Coupon payment * [1-(1+0.0736)^-22]/0.0736 + 1000/(1+0.0736)^22 = 800.26

=>

Coupon payment = 55

coupon rate = 55/1000 = 5.5%